OpenStudy (anonymous):

Find the valueof f(1)when f(x) = sin−1x+2tan−1x

OpenStudy (dumbcow):

sin^-1 (1) = pi/2 tan^-1 (1) = pi/4 --> f(1) = pi/2 + 2pi/4 = pi

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

your welcome